OpenStudy (anonymous):

Which of the following represents the vertical asymptotes of the function f(x) =(3x - 8)/(x^2 - 3x - 10)?

OpenStudy (anonymous):

x = 2 and x = 5 x = –2 and x = 5 x = 2 and x = –5 x = –2 and x = –5

OpenStudy (amistre64):

verts are what makes the denom go zero, but that do not have a canceling factor in the numerator

OpenStudy (amistre64):

it looks like the options are simple enough; just plug them into the denominator to see what zeros it out

OpenStudy (anonymous):

i dont understand how to do it though. I'm not good with asymptotes and I don't know what they mean when they say they have to be vertical.

OpenStudy (anonymous):

do I have to graph it? would that help?

OpenStudy (asnaseer):

do you know how to factorise the denominator?

OpenStudy (anonymous):

would it be (x - 5)(x + 2)?

OpenStudy (asnaseer):

yes

OpenStudy (asnaseer):

now see what valus of x would make this equal to zero

OpenStudy (asnaseer):

i.e. what values of x satisfy this equation:\[(x-5)(x+2)=0\]

OpenStudy (amistre64):

make sure theres no canceling factor in the numerator

OpenStudy (asnaseer):

yes ^^

OpenStudy (anonymous):

so is the answer x = -2 and x = 5 ?

OpenStudy (asnaseer):

what we have done so far is:\[\frac{3x-8}{x^2-3x-10}=\frac{3x-8}{(x-5)(x+2)}\]do you think there are any common factors between the numerator and denominator here?

OpenStudy (asnaseer):

yes - your answer is correct what @amistre64 was trying to show is that before the step of finding what values of x make (x-5)(x+2)=0, it is important to cancel out any common factors between the numerator and denominator.

OpenStudy (anonymous):

ohhh ok thanks

OpenStudy (asnaseer):

e.g., lets say your equation was actually:\[\frac{x+2}{(x-5)(x+2)}\]

OpenStudy (asnaseer):

then we would first cancel out the (x+2) terms

OpenStudy (asnaseer):

leaving 1/(x-5)

OpenStudy (asnaseer):

so the only asymptote here would be x=5